A population has a mean of 70 and a standard deviation of 6. A random sample of 15 measurements is drawn from this population.


Find the Value of μ and n

1
Expert's answer
2022-04-22T02:53:51-0400

"\mu_{\bar X}=\mu=70."

"\sigma_{\bar X}=\frac{\sigma}{\sqrt{n}}=\frac{6}{\sqrt{15}}=1.55."